# ArchRAG: Attributed Community-based Hierarchical Retrieval-Augmented Generation

### v1

@@ B.2. Implementation details
        {"description": "Description of point 1 [Data: Reports (report ids)]", "score": score_value},
        {"description": "Description of point 2 [Data: Reports (report ids)]", "score": score_value},
@@
Merge Prompt: # Role You are a helpful assistant responding to questions and may use the provided data as a reference. # Goal You should incorporate insights from all the reports from multiple analysts who focused on different parts of the dataset to support your answer. Please note that the provided information may contain inaccuracies or be unrelated. If the provided information does not address the question, please respond using what you know. namely, A response that utilizes the provided information, ensuring that all irrelevant details from the analysts’ reports are removed. A response to the user’s query based on your existing knowledge when ¡Analyst Reports¿ is empty. The final response should merge the relevant information into a comprehensive answer that clearly explains all key points and implications, tailored to the appropriate response length and format. Note that the analysts’ reports provided below are ranked in the descending order of importance. Do not include information where the supporting evidence for it is not provided. # Target response length and format
@@
Figure 10. The prompt for adaptive filtering-based generation.

### v3

@@ Appendix B Experimental details
        "description": "Description of point 1",
        "score": score_value
    // ... more points]}
@@
Merge Prompt: # Role You are a helpful assistant responding to questions and may use the provided data as a reference. # Goal You should incorporate insights from all the reports from multiple analysts who focused on different parts of the dataset to support your answer. Please note that the provided information may contain inaccuracies or be unrelated. If the provided information does not address the question, please respond using what you know: A response that utilizes the provided information, ensuring that all irrelevant details from the analysts’ reports are removed. A response to the user’s query based on your existing knowledge when ¡Analyst Reports¿ is empty. The final response should merge the relevant information into a comprehensive answer that clearly explains all key points and implications, tailored to the appropriate response length and format. Note that the analysts’ reports provided below are ranked in the descending order of importance. Do not include information where the supporting evidence for it is not provided. # Target response length and format
@@
Figure 19: The prompt for adaptive filtering-based generation.